2. One of the undersecretaries of the Department of Labor and Employment (DOLE) claims that the average salary of a civil engineer is Php 18,000. A sample of 19 civil engineer's salary has a mean of Php 17,350 and a standard deviation of Php 1,230. Is there enough evidence to reject the undersecretary's claim at a = 0.01?


Solution:

step 1: null and alternative hypothesis

H0:__________

H1:__________


step 2: Significance level a =__________


step 3: test statistic


step 4: decision rule


step 5: decision


1
Expert's answer
2022-06-08T14:52:11-0400

1. The following null and alternative hypotheses need to be tested:

"H_0:\mu=18000"

"H_1:\mu\not=18000"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=18" and the critical value for a two-tailed test is "t_c =2.87844."

The rejection region for this two-tailed test is "R = \{t:|t|>2.87844\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{17350-18000}{1230/\sqrt{19}}=-2.3035"


Since it is observed that "|t|=2.3035<2.87844=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for two-tailed, "df=18" degrees of freedom, "t=-2.3035" is "p=0.033391," and since "p=0.033391>0.01=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is different than 18000, at the "\alpha = 0.01" significance level.